OpenStudy (babystar16):

the length of a rectangle is 5 more than twice the width. the perminter is 130 in. what is the area?

OpenStudy (anonymous):

Let length be L. Let width by W. L = 2W + 5 -----(1) 2L + 2W = 130 L + W = 65 ------(2) Sub. (1) into (2) (2W + 5) + W = 65 3W + 5 = 65 3W = 60 W = 20 L = 2(20) + 5 = 45 Hence Area = 45 * 20 = 900

OpenStudy (babystar16):

so the area is 900?????

OpenStudy (anonymous):

yes

OpenStudy (babystar16):

thank u